# Simultaneous Gesture Classification and Localization with an Automatic Gesture Annotation Model

### v1

@@ 5 Experiments and Analysis
Table 1: Contents of the datasets. Static gestures are characterized by keeping a fixed hand pose for a minimum amount of time, while dynamic gestures are characterized by a single trajectory with an unchanged hand pose or with finger articulation over time. Fine dynamic gestures are characterized by a single trajectory with changing hand pose.
@@
Table 2: Ablation study results by adding our multiple design parts step by step. NNLE is our metric for gesture localization (Normalized Nucleus Labeling Error). The numbers in the brackets indicate std.
We conducted a series of experiments to evaluate our proposed annotation model.
5.1 Evaluation Datasets
In order to ensure a thorough evaluation of our Model, we perform our experiments on two publicly available datasets, SHREC’2021 [7] and Online DHG [8]. These datasets consist of sequences of unsegmented gesture data, encompassing various types of gestures. The specific details of the two datasets are as follows: Both datasets already provided a predefined split between training and testing data. Thus, we conducted evaluations on the designated testing sets.
5.2 Evaluation Metrics
@@
    Gesture Classification Accuracy: Gesture annotation would generate a sequence of classification output. We use the Levenshtein distance (also known as minimum edit distance) to evaluate the gesture classification performance. It is a metric defined as the minimum number of single-character (in our case, classification output of each frame) insertions, deletions, and substitutions required to transform one string into another. The accuracy for recognition performance is defined as:
    $Accuracy=1-\frac{\operatorname{levenshtein}(y_{predict},y_{true})}{\operatorname{length}(y_{true})}$ (1)
@@
    $NNLE=\frac{idx_{activation}-(idx_{start}+idx_{end})/2+1}{idx_{end}-idx_{start}+1}$ (2)
@@
5.3 Training Details
Our implementation was based on TensorFlow 2. We utilized the Adam optimizer with a starting learning rate of 0.0001. To prevent overfitting, we employed early stopping with a patience of 5. For both the Online DHG and SHREC’2021 datasets, we used a window length of 200 for our model. It is important to note that the length of an individual gesture within the datasets ranged from 20 to 50 frames, corresponding to a duration of approximately 0.4 to 1 second.
5.4 Ablation Study
@@
Table 2 illustrates the ablation study results. Compared to the baseline (A-ResLSTM trained with Cross-Entropy (CE) loss), our model significantly reduces the gesture localization error by 0.3 (71.4%) and 0.33 (75.0%). Regarding gesture classification accuracy, our model design achieves improvements of 4.3% on the SHREC’2021 dataset and 3.4% on the Online DHG dataset compared to the baseline.
@@
6 Conclusion
@@
This paper presents a model that enables accurate and automated annotation of gesture data, encompassing gesture classification and localization. The ablation study reveals that this novel annotation model design surpasses the baseline model, achieving a 4.3% higher class labeling accuracy and a 71.4% improvement in nucleus localization accuracy on the SHREC’2021 dataset (3.4% and 75.0% respectively on the Online DHG dataset).

### v2

@@ V Experiments and Analysis
TABLE I: Contents of the datasets. Static gestures are characterized by keeping a fixed hand pose for a minimum amount of time, while dynamic gestures are characterized by a single trajectory with an unchanged hand pose or with finger articulation over time. Fine dynamic gestures are characterized by a single trajectory with changing hand pose.
@@
TABLE II: Ablation study results by adding our multiple design parts step by step. NNLE is our metric for gesture localization (Normalized Nucleus Labeling Error). The numbers in the brackets indicate std.
We conducted a series of experiments to evaluate our proposed annotation model and the framework.
V-A Evaluation Datasets
In order to ensure a thorough evaluation of our framework, we perform our experiments on two publicly available datasets. These datasets consist of sequences of unsegmented gesture data, encompassing various types of gestures. The specific details of the two datasets are as follows:
  1. 1.
    SHREC’2021 [3]: The dataset utilized in this study consists of 18 distinct gesture classes belonging to various types. It comprises a total of 180 gesture sequences, each carefully designed to incorporate 3 to 5 gestures, accompanied by additional semi-random hand movements labeled as non-gestures. The original dictionary comprises 18 gestures, encompassing both static and dynamic gestures.
  2. 2.
    Online DHG [5]: The dataset used in this study comprises 14 distinct gesture classes from various categories. It consists of 280 sequences where each sequence contains 10 unsegmented gestures occurring sequentially. The data in this dataset is represented in the form of skeleton data, specifically capturing 22 joints for each hand skeleton. The skeleton data was collected using a Leap Motion device.
Both datasets already provided a predefined split between training and testing data. Thus, we conducted evaluations on the designated testing sets.
V-B Evaluation Metrics
@@
    Gesture Classification Accuracy: Gesture annotation would generate a sequence of classification output. We use the Levenshtein distance (also known as minimum edit distance) to evaluate the gesture classification performance [18]. It is a metric defined as the minimum number of single-character (in our case, classification output of each frame) insertions, deletions, and substitutions required to transform one string into another. The accuracy for recognition performance is defined as:
    $1-\frac{\operatorname{levenshtein}(y_{predict},y_{true})}{\operatorname{length}(y_{true})}$ (3)
@@
    $\frac{idx_{activation}-(idx_{start}+idx_{end})/2+1}{idx_{end}-idx_{start}+1}$ (4)
@@
V-C Training Details
Our implementation was based on TensorFlow 2. We utilized the Adam optimizer [17] with a starting learning rate of 0.0001. To prevent overfitting, we employed early stopping with a patience of 5. For both the Online DHG and SHREC’2021 datasets, we used a window length of 200 for our model. It is important to note that the length of an individual gesture within the datasets ranged from 20 to 50 frames, corresponding to a duration of approximately 0.4 to 1 second.
V-D Ablation Study
@@
Table II illustrates the ablation study results. Compared to the baseline (A-ResLSTM trained with Cross-Entropy (CE) loss), our model significantly reduces the gesture localization error by 0.30 (71.4%) and 0.33 (75.0%). Regarding gesture classification accuracy, our model design achieves improvements of 4.3% on the SHREC’2021 dataset and 3.4% on the Online DHG dataset compared to the baseline.
@@
V-E Annotation Framework Evaluation
SHREC’2021 [3] Online DHG [5]
Labeled Sequences $\sim 40$ $\sim 80$ $\sim 120$ $\sim 70$ $\sim 140$ $\sim 210$
Baseline [43] 30.0 ($\Delta$=22%) 69.0 ($\Delta$=11%) 83.3 ($\Delta$=6%) 34.9 ($\Delta$=25%) 68.3 ($\Delta$=10%) 81.6 ($\Delta$=5%)
Our Annotation Framework 26.0 ($\Delta$=18%) 73.0 ($\Delta$=15%) 88.3 ($\Delta$=11%) 28.9 ($\Delta$=19%) 75.3 ($\Delta$=17%) 89.6 ($\Delta$=13%)
TABLE III: Evaluation of the effectiveness of the pseudo-labeled dataset produced from our proposed annotation framework. The numbers indicate the gesture classification accuracy of the downstream gesture recognition model fine-tuned from the pseudo-labeled dataset. The $\Delta$ indicates the accuracy improvement between the pre-trained accuracy and the fine-tuned accuracy. The second row indicates the size of the labeled subset. These labeled sequences are used for pre-training the downstream gesture recognition model and the annotation model. Please refer to Figure 1 for a detailed illustration of the workflow of the annotation framework. Note that SHREC’2021 dataset has a total of 180 sequences, and Online DHG dataset has a total of 280 sequences.
To evaluate the effectiveness of our framework, we incorporate a similar pseudo-labeling process into the baseline model described in Section V-D. This modified version serves as a baseline framework for comparison.
V-E1 Baseline
It is important to note that the pseudo labeling semi-supervised learning pipeline we developed specifically for our proposed annotation model cannot be directly applied to the baseline model, as the baseline model is trained using Cross-Entropy (CE) loss. Therefore, we made modifications to adapt the pseudo-labeling process to the baseline model, enabling a fair comparison between the two frameworks.
Formally, we denote the CE loss function between two probability distributions $p$ and $q$ as $H(p,q)$. Let $q_{b}=p(\textbf{y}\mid\alpha(\textbf{u}_{b}))$ be the class-conditional probability distribution produced by the model given a weakly-augmented version of a given unlabeled sample: u. Then we use $\hat{q}_{b}=\operatorname*{arg\,max}(q_{b})$ through the path decoding algorithm as a pseudo-label, and we enforce the CE loss against the model’s output for a strongly-augmented version of $\textbf{u}_{b}$:
$L=\frac{1}{\mu B}\sum_{b=1}^{\mu B}\mathbb{I}(\operatorname*{arg\,max}(q_{b})\geq\tau)H(\hat{q}_{b}\mid,p(\textbf{y}\mid\mathrm{A}(\textbf{u}_{b})))$
where $\tau$ is the threshold. The rest of the semi-supervised learning pipeline stays the same as in Section IV-C.
V-E2 Procedure
The ultimate objective of our framework is to generate high-quality pseudo-labels that can effectively fine-tune and enhance downstream gesture recognition models. To evaluate the efficacy of our pipeline, we measure the performance improvement achieved by fine-tuning the initial pre-trained gesture recognition model using our generated pseudo-labels.
In the experiments, we create a subset of the training dataset that initially contains complete labels. From this subset, we remove the labels and leverage our framework to generate pseudo-labels for this particular portion of the data. The size of the remaining subset of labeled data becomes the adjustable parameter in our experiments. Now, we have two subsets, one labeled subset and one unlabeled subset. This labeled subset is utilized for pre-training both the annotation model and the downstream gesture recognition model, and the unlabeled subset is utilized in the pseudo labeling-based semi-supervised learning pipeline. Such a process is demonstrated in Figure 1. We use one of the state-of-the-art gesture recognition models, ST-GCN, from [3] as the downstream gesture recognition model. The performance of the framework directly impacts the quality of the pseudo-labels produced from the annotation framework. The quality of the pseudo-labels then impacts the improved accuracy and final accuracy of the final downstream recognition model which is fine-tuned from the pseudo-labels. This is explained in Figure 1. We use gesture classification accuracy in Equation 3 to measure the accuracy of the downstream gesture recognition model.
V-E3 Results
The performance and improvements of the fine-tuned downstream recognition model are summarized in Table III. It is important to note that these results pertain to the downstream model and are not directly comparable to those in Table II.
Through the semi-supervised learning pipeline, both our new annotation framework and the baseline framework demonstrate significant improvements in the performance of the downstream model, as indicated by the $\Delta$ values, from 5% to 25%. The magnitude of improvement is directly influenced by the size of the unlabeled dataset, with larger datasets yielding greater improvements through this pipeline.
Furthermore, our evaluation results highlight that the pseudo-labeled data generated by our framework generally outperforms the pseudo-labeled data generated by the baseline framework. The quality of the pseudo labels is reflected in both the improved accuracy and the final accuracy of the downstream model. It is noteworthy that the baseline framework initially outperforms our new framework when using a small set of labeled data. This can be attributed to the higher number of trainable parameters in the output layer of our annotation model, which employs a many2many design and faces challenges in achieving convergence with the CTC loss. However, as the size of the labeled set increases, our new framework surpasses the baseline.
The superior performance of our framework can be attributed to the annotation model’s ability to train on pseudo-labeled data that may not be accurately segmented initially in the semi-supervised learning loop (pipeline). This is in contrast to the baseline framework, whose model heavily relies on accurately segmented data when trained with CE loss. These results highlight the potential of our annotation model and the semi-supervised pipeline to significantly enhance downstream gesture recognition models, showcasing the advantages of our approach.
VI Limitation and Future Work
It is important to acknowledge several limitations in our work. Firstly, although our proposed model exhibits superior performance in gesture data annotation for class labeling and nucleus localization, we must acknowledge that the training process of our new model is more time-consuming compared to the baseline. While training efficiency is not the primary focus since annotation can occur offline, there is room for future improvement in this aspect. Furthermore, our current framework still relies on the availability of labeled data to initiate the process. We have not explored a fully unsupervised version in this study, which presents an intriguing avenue for future research.
VII Conclusion
@@
This paper presents a framework that enables accurate and automated annotation of gesture data, encompassing gesture classification and localization. The framework consists of two key components: (1) a novel annotation model that optimizes both gesture classification and localization concurrently, and (2) a semi-supervised learning pipeline incorporating pseudo-labeling. The ablation study reveals that this novel annotation model design surpasses the baseline model, achieving a 4.3% higher class labeling accuracy and a 71.4% improvement in nucleus localization accuracy on the SHREC’2021 dataset (3.4% and 75.0% respectively on the Online DHG dataset). Moreover, we also demonstrate that the pseudo-labeled data generated by our framework significantly enhances the performance of a pre-trained downstream gesture recognition model through fine-tuning, resulting in improvements ranging from 11% to 18% on the SHREC’2021 and Online DHG datasets.